OpenStudy (anonymous):

Use basic identities to simplify the expression. cos θ - cos θ sin2θ

OpenStudy (anonymous):

Well to begin do you know the double angle identity for sin?

OpenStudy (anonymous):

sin^2x + cos^2x = 1

OpenStudy (anonymous):

Not quite that is one of the Pythagorean Identities. The double angle identity for sin is: \[\sin2\theta=2\sin \theta \cos \theta\]

OpenStudy (anonymous):

does that help at all?

OpenStudy (anonymous):

a little, but I'm still confused. Wouldn't that give me cosx - sin^2x?

OpenStudy (anonymous):

wait is that 2 in front of the theta an exponent or a 2 being multiplied by theta?

OpenStudy (anonymous):

It's an exponent

OpenStudy (anonymous):

Does this question have choices at all?

OpenStudy (anonymous):

sec2θ sin θ tan2θ cos3θ

OpenStudy (anonymous):

well first if this does help at all look for something in common, what can you take out of the problem or factor out that simplifies the expression. Think of something like a-ab=a(1-b)

OpenStudy (anonymous):

ummmmm.... can you factor out cos?

OpenStudy (anonymous):

yes! and your left with cosθ(1−sin2θ) Now think of the Pythagorean identity that you mentioned before, what can \[1-\sin^2\theta\]\ turn into?

OpenStudy (anonymous):

cos^2 x! So the answer would be cos^3 x!

OpenStudy (anonymous):

yes! correct!

OpenStudy (anonymous):

or as your choice is \[\cos^3\theta\]

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

Your welcome :)